OpenStudy (anonymous):

H(t)=-16t2+96t Find the vertex. (show steps please, I'm so confused!)

OpenStudy (anonymous):

vertex use \[-\frac{b}{2a}\]

OpenStudy (anonymous):

in this case a = -16, b = 96 to \[-\frac{b}{2a}=-\frac{96}{2\times -16}=3\]

OpenStudy (anonymous):

3 is the first coordinate and \[H(3)=-16\times 3^2+96\times t =144\]

OpenStudy (anonymous):

so vertex is (3,144)

OpenStudy (anonymous):

if the question is "how high does the ball go?" the answer is the max is 144 ft after 3 seconds

OpenStudy (anonymous):

steps clear?